### v1

@@ Abstract
Reconstructing a sequence of sharp images from the blurry input is crucial for enhancing our insights into the captured scene and poses a significant challenge due to the limited temporal features embedded in the image. Spike cameras, sampling at rates up to 40,000 Hz, have proven effective in capturing motion features and beneficial for solving this ill-posed problem. Nonetheless, existing methods fall into the supervised learning paradigm, which suffers from notable performance degradation when applied to real-world scenarios that diverge from the synthetic training data domain. Moreover, the quality of reconstructed images is capped by the generated images based on motion analysis interpolation, which inherently differs from the actual scene, affecting the generalization ability of these methods in real high-speed scenarios. To address these challenges, we propose the first self-supervised framework for the task of spike-guided motion deblurring. Our approach begins with the formulation of a spike-guided deblurring model that explores the theoretical relationships among spike streams, blurry images, and their corresponding sharp sequences. We subsequently develop a self-supervised cascaded framework to alleviate the issues of spike noise and spatial-resolution mismatching encountered in the deblurring model. With knowledge distillation and re-blurring loss, we further design a lightweight deblur network to generate high-quality sequences with brightness and texture consistency with the original input. Quantitative and qualitative experiments conducted on our real-world and synthetic datasets with spikes validate the superior generalization of the proposed framework. Our code, data and trained models will be available at https://github.com/chenkang455/S-SDM.
@@
Keywords:
Motion Deblur Spike Camera Self-supervised
Figure 1: Illustration of the superiority of our self-supervised framework (S-SDM) over supervised methods, where “Refer to caption” symbolizes active network training and “Refer to caption” indicates that network parameters remain unchanged during this stage. Supervised methods, while effective on synthetic datasets, suffer from a significant performance decline when applied to real-world datasets, primarily due to data distribution discrepancies. In contrast, our self-supervised framework, necessitating no Ground Truth (GT) for training, seamlessly bridges this dataset gap through fine-tuning on real-world datasets.
@@ 1 Introduction
Traditional cameras, constrained by their exposure-based imaging mechanism, often produce blurry images when capturing fast-moving objects or during camera movement throughout the exposure process [19, 28]. While these blurry images lose significant details, the ability to recover dynamic motion trajectories from the static blurry input becomes critically important. However, the inherent challenge lies in the limited motion features available within blurry frames, leading to potential ambiguities such as multiple motion trajectories corresponding to the same blurry input. This is exemplified by scenarios where two objects move along the same trajectory but in opposite directions [22, 23, 30, 31, 25], rendering the task of motion deblurring ill-posed. Recent advancements in learning-based approaches [32, 12, 38] seek to address this challenge by establishing direct mappings from blurry inputs to sharp sequences in a self-supervised manner. Despite these efforts, traditional cameras struggle to capture fine details in high-speed motion due to their exposure constraints, thus limiting the effectiveness of these methods in scenarios not covered by the training datasets.
@@
In recent years, neuromorphic cameras[9, 10], leveraging their ultra-high temporal resolution and high dynamic range, have found widespread use in many fields, including computational imaging, computer vision, and robotics. These cameras, including event and spike cameras, are distinguished by their ability to produce high temporal resolution outputs directly tied to changes in light intensity. Specifically, event cameras generate events in areas where light intensity changes [4], while spike cameras capture the absolute brightness of the scene at each pixel, offering a stream of spikes as output [10]. This distinctive feature endows spike cameras with a significant advantage in capturing and recovering texture information from scenes with rapid motion as validated by numerous recent studies [39, 37, 40, 5, 33, 6].
@@
Recent studies [2, 8] have explored the potential of RGB-Spike fusion, i.e., harnessing the strengths of both traditional and spike cameras to reconstruct sharp sequences from blurry inputs. However, their frameworks are constrained within the supervised learning paradigm, which necessitates extensive datasets comprising pairs of blurry and sharp images, as well as spike sequences. While synthetically acquiring such paired data, as demonstrated in previous studies [8, 4, 23], is feasible, collecting them in real-world scenarios presents several challenges: (1) high-speed cameras are prohibitively expensive and not readily deployable in many settings; (2) spatial-temporal calibration between spike cameras and high-speed RGB cameras complicates the data collection process. These challenges render the fine-tuning of supervised methods on real-world datasets impossible, further leading to their performance deterioration in such environments as shown in Fig. 1. The resulting degradation in image quality, manifesting as color distortion, brightness inconsistency, and inaccurate texture restoration, is mainly caused by the disparity between synthetic and real-world datasets, especially in terms of the density of spike stream, spike generation mechanism, and blurry image generation. Moreover, the effectiveness of supervised methods is inherently limited by the ground truth sequences created through motion analysis interpolation algorithms [21, 11], which inherently differs from the real-world scene and thus affects the model’s generalization ability.
@@
To overcome these issues, we propose the first-of-its-kind Self-supervised Spike-guided Deblurring Model (S-SDM), capable of recovering the continuous sharp sequence from a single blurry input with the assistance of low-resolution spike streams. We begin with a theoretical analysis of the relationship between spike streams, blurry images, and sharp sequences, leading to the development of our Spike-guided Deblurring Model (SDM). We further construct a self-supervised processing pipeline by cascading the denoising network and the super-resolution network to reduce the sensitivity of the SDM to spike noise and its reliance on spatial-resolution matching between the two modalities. To reduce the time consumption and enhance the utilization of spatial-temporal spike information within this pipeline, we further design a Lightweight Deblurring Network (LDN) and train it based on pseudo-labels from the teacher model, i.e., the established self-supervised processing pipeline. By further introducing re-blurring loss during LDN training, we achieve better restoration performance and faster processing speed compared to the processing-lengthy and structure-complicated teacher model. To validate the performance of our S-SDM across various scenarios, we build an RGB-Spike binocular system and propose the first spatially-temporally calibrated real-world spike blur dataset in this community. Quantitative and qualitative experiments conducted on our real-world and synthetic datasets with spikes validate the superior generalization of our S-SDM.
In summary, the key contributions of this paper are:
@@
    We develop a self-supervised spike-guided image deblurring framework, addressing the performance degradation due to the synthetic-real domain gap that existed in previous supervised methods.
@@
    We propose a real-world blur dataset and experiments on real-world and synthetic datasets validate the superior generalization of our S-SDM.
@@ 2 Related Work
Spike Camera. The spike camera, inspired by the primate retina, stands apart from conventional cameras with its ability to generate synchronous spike streams for each pixel at extremely low latency. This distinct feature provides significant advantages in various applications such as high-speed imaging [39, 37, 40, 5, 33, 6], optical flow estimation [35], object detection [36], depth estimation [27], motion deblurring [8], and occlusion removal [26].
@@ 3.1 Preliminaries
Spike Camera Mechanism. Consider $\mathbf{L}(t)$ to represent the latent sharp frame at time $t$. Each pixel $p$ in the spike camera [10] has an integrator that accumulates the incoming photons at a high frequency. Once the cumulative intensity exceeds a predefined threshold $C$ at time $t_{e}$, pixel $p$ emits a spike, and the accumulation of photons is reset to zero. This process can be mathematically described as follows:
@@
where $t_{s}$ denotes the firing time of the previous spike. While the spike camera is capable of generating asynchronous spike streams akin to that of event cameras, its effectiveness is constrained by the inherent limitations of its physical circuitry, which necessitates reading spikes at a predetermined sampling rate. We denote the generated spike stream as $\mathcal{S}\in\{0,1\}^{K\times H\times W}$, where $H$ and $W$ signify the height and width of the image, and $K$ represents the length of the spike sequence.
@@ 3.2 Theoretical Analysis
where $\mathbf{B}_{g}$ denotes the grayscale version of the blurry input, and $N_{\mathcal{T}}$ denotes the total number of spikes accumulated over the exposure period, as shown in Fig. 2(a). The accumulation $N_{\mathcal{T}}$ is calculated as $N_{\mathcal{T}}=\sum_{i=1}^{K}\mathcal{S}[i]$, with $\mathcal{S}[i]$ indicating the $i$-th frame of the spike stream.
Within the exposure $\mathcal{T}$, we consider a shorter spike sequence centered around the $t$ moment $S_{\mathcal{T^{\prime}}}\in\{0,1\}^{K^{\prime}\times 1\times H\times W}$, satisfying $K^{\prime}\ll K$, $t\in\mathcal{T}$, and $\mathcal{T^{\prime}}\subset\mathcal{T}$. Similar to Eq. 5, we can derive the relationship between the short-exposure gray image $B_{g}(t,\mathcal{T}^{\prime})$ and the short spike stream $\mathcal{S}_{\mathcal{T^{\prime}}}$ as follows:
@@
where $T^{\prime}\ll T$ represents the short exposure time. We adopt a notation in which the bolded symbol $\mathbf{B}_{g}$ signifies the image captured by conventional cameras, whereas the non-bolded expression $B_{g}(t,\mathcal{T}^{\prime})$ corresponds to images reconstructed from the spike stream, as depicted in Fig. 2(a).
Given the observation that the color information of adjacent pixels often exhibits similarity under the premise of minor motion amplitude, we postulate that the colors of the blurry input $\mathbf{B}$ and the short-exposure image $B(t,\mathcal{T}^{\prime})$ are identical. This assumption implies that the intensity proportion among RGB channels in two images $\boldsymbol{\alpha}_{c}/\alpha_{c}(t,\mathcal{T}^{\prime})$ is approximately equivalent, i.e., $\boldsymbol{\alpha}_{c}\approx\alpha_{c}(t,\mathcal{T}^{\prime})$, satisfying $\boldsymbol{\alpha}_{c}=\mathbf{B}_{g}/\mathbf{B}_{c}$ and $\alpha_{c}(t,\mathcal{T}^{\prime})=B_{g}(t,\mathcal{T}^{\prime})/B_{c}(t,\mathcal{T}^{\prime})$. More details can be found in the supplementary materials.
Upon establishing it, we move forward to build a mathematical relation between the two images. By substituting the gray channel $g$ with color channel $c$ and dividing Eq. 5 by Eq. 6, we efficiently eliminate the unknown threshold $C$ and weights, leading to the following equation:
$B_{c}(t,\mathcal{T}^{\prime})=\mathbf{B}_{c}\cdot\frac{N_{\mathcal{T}^{\prime}}}{N_{\mathcal{T}}}\cdot\frac{T}{T^{\prime}}.$ (7)
By applying Eq. 7 to RGB three channels, we explicitly establish the relationship between the color blurry input $\mathbf{B}$, the color short-exposure image $B(t,\mathcal{T}^{\prime})$ and the spike stream $S_{\mathcal{T}}$ as shown in Fig. 2(a). Since the exposure period $\mathcal{T^{\prime}}$ is relatively short, it is reasonable to assume that the scene remains static. In this context, we interpret the short-exposure image $B(t,\mathcal{T}^{\prime})$ as the latent sharp frame $\mathbf{L}(t)$, allowing us to modify Eq. 7 as follows:
@@
To this end, we have conducted a comprehensive theoretical analysis of the spike-guided motion deblurring task, which is neglected in prior learning-based motion deblur methodologies [2, 8]. For further discussion readability, we refer to Eq. 8 as the Spike-guided Deblurring Model (SDM) as shown in Fig. 2(a), which is analogous to the baseline motion deblur model EDI [18] in event camera.
Figure 2: The schematic diagram presents the SDM model (a), the denoising network (b), and the SR network (c). In these diagrams, “Refer to caption” symbolizes the network is trained during this stage. Additionally, the black hole in (b) signifies that the information in this area is not accessible to the BSN.
@@ 3.3 Self-supervised Spike-guided Deblurring Model
    The deblurred image $B(t,\mathcal{T}^{\prime})$ suffers from noise-related degradation attributed to the lack of adequate spike information during the short exposure $\mathcal{T}^{\prime}$.
@@
To tackle these limitations, we first devise a lengthy yet feasible processing pipeline based on the SDM, as illustrated in the Fig. 3. Initially, we utilize Eq. 6 to convert the short-exposure spike stream into an image, followed by a sequential cascade of the Denoising Network and the Super-Resolution Network, which are tasked with denoising and resolution-enhancement respectively in a self-supervised manner. Subsequently, the deblurred image is obtained through the SDM, with inputs comprising both the blurry image and the enhanced short-exposure image. In the following, we delve into the details of the Denoising Network and Super-Resolution Network.
Denoising Network. To this end, We leverage the powerful self-supervised denoising capabilities of the Blind Spot Network (BSN) [13, 1, 24, 5, 14, 7]. The core idea of BSN involves designing the blind-spot mechanism, which compels the network to estimate the value of each pixel based on the intensities of its surrounding pixels, as illustrated in Fig. 2(b). Under the premise that the noise is random and independent, the BSN can be trained to accurately predict the value of clean pixels and deduce the appropriate values for noisy pixels from their surroundings.
In the pre-training part of our processing pipeline, the short-exposure noisy image $B(t,\mathcal{T}^{\prime})$ serves as both the input and the supervision signal for the BSN [13] with the loss function formulated as:
$\mathcal{L}_{BSN}=||B(t,\mathcal{T}^{\prime})-\text{BSN}(B(t,\mathcal{T}^{\prime});\Theta_{1})||_{2}^{2},$ (9)
where $\Theta_{1}$ represents the parameters of the BSN. Once the BSN is trained, we feed the noisy image into it to yield the denoised image $\mathbf{L}_{bsn}(t)$.
Super-Resolution Network. In this task, we observe that the blurry input $\mathbf{B}$ from the RGB camera and the long-exposure image $B(f,\mathcal{T})$ from the spike camera, where $f$ is the central moment during the exposure shown in Fig. 2(a), exhibit similar characteristics. This motivates us to train the SR network based on these pairs as shown in Fig. 2(c) and subsequently apply it to enhance the resolution of the short-exposure image.
We leverage the well-explored network Enhanced Deep Super-Resolution network (EDSR) [15] as the backbone of our SR network, with the long-exposure image $B_{g}(f,\mathcal{T})\in\mathbb{R}^{1\times\frac{H}{4}\times\frac{W}{4}}$ as the input and the blurry gray input $\mathbf{B}_{g}\in\mathbb{R}^{1\times H\times W}$ as supervision. The loss function is formulated as follows:
$\mathcal{L}_{SR}=||\mathbf{B}_{g}-\text{EDSR}(B_{g}(f,\mathcal{T});\Theta_{2})||_{2}^{2},$ (10)
where $\Theta_{2}$ represents the parameters of the EDSR. Subsequently, we freeze the network parameters and apply it to the denoised short-exposure image $\mathbf{L}_{bsn}(t)$, yielding the resolution-enhanced image $\mathbf{L}_{bsn}^{\uparrow}(t)$.
Figure 3: The schematic diagram of our proposed distillation self-supervised framework, where “Recon” denotes the short-exposure image reconstruction corresponding to Eq. 6. In this diagram, the “Refer to caption” symbol denotes that the network parameters are frozen during this stage, while the “Refer to caption” indicates that certain computations are executed in a non-network manner.
Lightweight Deblur Network. While the aforementioned processing pipeline successfully achieves the multi-modality fusion of the High-spatial/Low-temporal blurry input and the High-temporal/Low-spatial spike stream, several aspects still need refinement:

### v4

@@ Abstract
Reconstructing a sequence of sharp images from the blurry input is crucial for enhancing our insights into the captured scene and poses a significant challenge due to the limited temporal features embedded in the blurry image. Spike cameras, sampling at rates up to 40,000 Hz, have proven effective in capturing motion features and beneficial for solving this ill-posed problem. Nonetheless, existing methods fall into the supervised learning paradigm, which suffers from notable performance degradation when applied to real-world scenarios that diverge from the synthetic training data domain. To address this challenge, we propose the first self-supervised framework for the task of spike-guided motion deblurring. Our approach begins with the formulation of a spike-guided deblurring model that explores the theoretical relationships among spike streams, blurry images, and their corresponding sharp sequences. We subsequently develop a self-supervised cascaded framework to alleviate the issues of spike noise and spatial-resolution mismatching encountered in the deblurring model. With knowledge distillation and reblur loss, we further design a lightweight deblur network to restore high-quality sequences with brightness and texture consistency with the original input. Quantitative and qualitative experiments conducted on our real-world and synthetic datasets with spikes validate the superior generalization of the proposed framework. Our code, data and trained models are available at https://github.com/chenkang455/S-SDM.
@@
Figure 1: Illustration of the superiority of our self-supervised framework (S-SDM) over supervised methods. Supervised methods, while effective on synthetic datasets, suffer from a significant performance decline when applied to real-world datasets, primarily due to data distribution discrepancies. In contrast, our self-supervised framework, necessitating no Ground Truth (GT) for training, seamlessly bridges this dataset gap through fine-tuning on real-world datasets.
@@ 1 Introduction
Traditional cameras, constrained by their exposure-based imaging mechanism, often produce blurry images when capturing fast-moving objects or during camera movement throughout the exposure process [20, 33]. While these blurry images lose significant details, the ability to recover dynamic motion trajectories from the static blurry input becomes critically important. However, the inherent challenge lies in the limited motion features available within blurry frames, leading to potential ambiguities such as multiple motion trajectories corresponding to the same blurry input. This is exemplified by scenarios where two objects move along the same trajectory but in opposite directions [24, 25, 36], rendering the task of motion deblurring ill-posed. Recent advancements in learning-based approaches [39, 13, 49] seek to address this challenge by establishing direct mappings from blurry inputs to sharp sequences in a supervised learning manner. Despite these efforts, traditional cameras struggle to capture fine details in high-speed motion due to their exposure constraints, thus limiting the effectiveness of these methods in scenarios not covered by the training datasets.
@@
In recent years, neuromorphic cameras [9, 11], leveraging their ultra-high temporal resolution and high dynamic range, have found widespread use in many fields, including computer vision and robotics. These cameras, including event and spike cameras, are distinguished by their ability to produce high temporal resolution outputs directly tied to changes in light intensity. Specifically, event cameras generate events in areas where light intensity changes [4], while spike cameras capture the absolute brightness of the scene at each pixel, offering a stream of spikes as output [11]. This distinctive feature endows spike cameras with a significant advantage [50, 48, 51, 5, 40, 6] in capturing and recovering sharp texture from scenes with rapid motion.
@@
Recent studies [2, 8] have explored the potential of RGB-Spike fusion, i.e., harnessing the strengths of both traditional and spike cameras to reconstruct sharp sequences from blurry inputs. However, their frameworks are constrained within the supervised learning paradigm, which necessitates extensive datasets comprising pairs of blurry and sharp images, as well as spike sequences. While synthetically acquiring such paired data, as demonstrated in previous studies [8, 4, 25], is feasible, collecting them in real-world scenarios presents the following challenges: (1) high-speed cameras are prohibitively expensive and not readily deployable in many settings; (2) spatial-temporal calibration between spike cameras and high-speed RGB cameras complicates the data collection process. These problems render the fine-tuning of supervised methods on real-world datasets challenging, further leading to their performance deterioration in such environments as shown in Fig. 1. The resulting degradation in image quality, manifesting as color distortion, brightness inconsistency, and inaccurate texture restoration, is mainly caused by the disparity between synthetic and real-world datasets, especially in terms of the density of spike stream, spike generation mechanism, and blurry image generation. Moreover, the effectiveness of supervised methods is inherently limited by the ground truth sequences created through motion analysis interpolation algorithms [23, 12], which inherently differs from the real-world scene and thus affects the model’s generalization ability.
@@
To overcome these issues, we propose the first-of-its-kind Self-supervised Spike-guided Deblurring Model (S-SDM), capable of recovering the continuous sharp sequence from a single blurry input with the assistance of low-resolution spike streams. We begin with a theoretical analysis of the relationship between spike streams, blurry images, and sharp sequences, leading to the development of our Spike-guided Deblurring Model (SDM). We further construct a self-supervised processing pipeline by cascading the denoising network and the super-resolution network to reduce the sensitivity of the SDM to spike noise and its reliance on spatial-resolution matching between the two modalities. To reduce the computational cost and enhance the utilization of spatial-temporal spike information within this pipeline, we further design a Lightweight Deblurring Network (LDN) and train it based on pseudo-labels from the teacher model, i.e., the established self-supervised processing pipeline. Further introducing reblur loss during LDN training, we achieve better restoration performance and faster processing speed than the processing-lengthy and structure-complicated teacher model. To validate the performance of our S-SDM across various scenarios, we build an RGB-Spike binocular system and propose the first spatially-temporally calibrated Real-world Spike Blur (RSB) dataset in this community. Quantitative and qualitative experiments conducted on the real-world and synthetic datasets validate the superiority of our method. In summary, our key contributions are:
@@
    We develop a self-supervised spike-guided image deblurring framework, addressing the performance degradation due to the synthetic-real domain gap in supervised methods.
@@
    We propose a real-world dataset RSB and experiments on GOPRO and RSB datasets validate the superior generalization of our S-SDM.
@@ 2 Related Work
Spike Camera. The spike camera [42, 47, 46, 44, 38], inspired by the primate retina, stands apart from conventional cameras with its ability to generate synchronous spike streams for each pixel at extremely low latency. This distinct feature provides significant advantages in various applications such as high-speed imaging [50, 48, 51, 5, 40, 6], optical flow estimation [43], object detection [45], depth estimation [32], motion deblurring [8], and occlusion removal [31].
@@ 3.1 Preliminaries
Spike Camera Mechanism. Consider $\mathbf{L}(t)$ to represent the latent sharp frame at time $t$. Each pixel $p$ in the spike camera [11, 10] has an integrator that accumulates the incoming photons at a high frequency. Once the cumulative intensity exceeds a predefined threshold $C$ at time $t_{e}$, pixel $p$ emits a spike, and the accumulation of photons is reset to zero. This process can be mathematically described as follows:
@@
where $t_{s}$ denotes the firing time of the previous spike. While the spike camera is capable of generating asynchronous spike streams akin to that of event cameras, its effectiveness is constrained by the inherent limitations of its physical circuitry, which necessitates reading spikes at a predetermined sampling rate. We denote the generated spike stream as $\mathcal{S}\in\{0,1\}^{K\times 1\times H\times W}$, where $H$ and $W$ signify the height and width of the image, and $K$ represents the length of the spike sequence.
@@ 3.2 Theoretical Analysis
where $\mathbf{B}_{g}$ denotes the grayscale version of the blurry input, and $N_{\mathcal{T}}$ denotes the total number of spikes accumulated over the exposure period. The accumulation $N_{\mathcal{T}}$ is calculated as $N_{\mathcal{T}}=\sum_{i=1}^{K}\mathcal{S}[i]$, with $\mathcal{S}[i]$ indicating the $i$-th frame of the spike stream.
Within the exposure $\mathcal{T}$, we consider a shorter spike sequence centered around the $t$ moment $\mathcal{S}_{\mathcal{T^{\prime}}}\in\{0,1\}^{K^{\prime}\times 1\times H\times W}$, satisfying $K^{\prime}\ll K$, $t\in\mathcal{T}$, and $\mathcal{T^{\prime}}\subset\mathcal{T}$. Similar to Eq. 5, we can derive the relationship between the short-exposure gray image $\mathbf{E}_{g}(t,\mathcal{T}^{\prime})$ and the short spike stream $\mathcal{S}_{\mathcal{T^{\prime}}}$ as follows:
@@
where $T^{\prime}\ll T$ represents the short exposure period.
Given the observation that the color information of adjacent pixels often exhibits similarity under the premise of minor motion amplitude, we postulate that the colors of the blurry input $\mathbf{B}$ and the short-exposure image $\mathbf{E}(t,\mathcal{T}^{\prime})$ are identical. This assumption implies that the intensity proportion among RGB channels in the blurry input $\alpha_{c}^{\mathbf{B}}$ and the short-exposure image $\alpha_{c}^{\mathbf{E}}(t,\mathcal{T}^{\prime})$ is approximately equivalent, satisfying $\alpha_{c}^{\mathbf{B}}=\mathbf{B}_{g}/\mathbf{B}_{c}$ and $\alpha_{c}^{\mathbf{E}}(t,\mathcal{T}^{\prime})=\mathbf{E}_{g}(t,\mathcal{T}^{\prime})/\mathbf{E}_{c}(t,\mathcal{T}^{\prime})$. More details can be found in the supplementary materials.
Upon establishing it, we move forward to build a mathematical relation between the blurry image and the spike stream. By substituting the gray channel $g$ with color channel $c$ and dividing Eq. 5 by Eq. 6, we efficiently eliminate the unknown threshold $C$ and weights $\alpha_{c}^{\mathbf{B}}/\alpha_{c}^{\mathbf{E}}(t,\mathcal{T}^{\prime})$, leading to the following equation:
$\mathbf{E}_{c}(t,\mathcal{T}^{\prime})=\mathbf{B}_{c}\cdot\frac{N_{\mathcal{T}^{\prime}}}{N_{\mathcal{T}}}\cdot\frac{T}{T^{\prime}}.$ (7)
By applying Eq. 7 to RGB three channels, we explicitly establish the relationship between the color blurry input $\mathbf{B}$, the color short-exposure image $\mathbf{E}(t,\mathcal{T}^{\prime})$ and the spike stream $\mathcal{S}_{\mathcal{T}}$ as shown in Fig. 19. Since the exposure period $\mathcal{T^{\prime}}$ is relatively short, it is reasonable to assume that the scene remains static. In this context, we interpret the short-exposure image $\mathbf{E}(t,\mathcal{T}^{\prime})$ as the latent sharp frame $\mathbf{L}(t)$, allowing us to modify Eq. 7 as follows:
@@
To this end, we have conducted a comprehensive theoretical analysis of the spike-guided motion deblurring task, which is neglected in prior learning-based motion deblur methodologies [2, 8]. For further discussion readability, we refer to Eq. 8 as the Spike-guided Deblurring Model (SDM), which is analogous to the baseline motion deblur model EDI [19] in event camera.
Figure 2: The schematic diagram of our proposed distillation self-supervised framework. The “Refer to caption" indicates that certain computations are executed in a non-network manner.
@@ 3.3.1 Processing Pipeline
    The deblurred image $\mathbf{E}(t,\mathcal{T}^{\prime})$ suffers from noise-related degradation due to the lack of adequate spike information during the short exposure $\mathcal{T}^{\prime}$.
@@
To overcome these limitations, we further cascade the self-supervised denoising network to eliminate the spike noise in $N_{\mathcal{T}^{\prime}}$ and super-resolution network to match spatial resolutions of the blurry image and spike stream, with the processing pipeline illustrated in the bottom of Fig. 2.
Denoising Network. We leverage the Blind Spot Network (BSN) [14, 1, 28, 5, 15, 7] to predict the clean spike accumulation $N_{\mathcal{T}^{\prime}}$ from the input short-exposure spike stream $\mathcal{S}_{\mathcal{T}^{\prime}}$. The core idea of BSN is to design the blind-spot strategy that compels the convolutional layer to estimate the clean value of each pixel solely based on its surrounding pixels.
Under the premise that the spike stochastic thermal noise is independent identically distributed [41], the BSN is trained to deduce sharp spike frames from the input, with the loss function formulated as:
$\mathcal{L}_{\text{BSN}}=||\text{BSN}(\mathcal{S}_{\mathcal{T}^{\prime}};\Theta_{\text{1}})-N_{\mathcal{T}^{\prime}})||_{2}^{2},$ (9)
where the denoised spike frame $\text{BSN}(\mathcal{S}_{\mathcal{T}^{\prime}};\Theta_{\text{1}})$ is denoted by $\mathbf{L}_{\text{bsn}}(t)$ for further analysis.
Super-Resolution Network. In this task, we observe that the blurry input $\mathbf{B}_{g}$ and the long-exposure spike frame $N_{\mathcal{T}}$ exhibit the same texture features as shown in Eq. 5. This observation motivates us to train the Super-Resolution (SR) network based on pairs of the blurry images and the long-exposure spike frames.
We leverage the well-explored Enhanced Deep Super-Resolution network (EDSR) [16] as the backbone of our SR network, with the loss function formulated as follows:
$\mathcal{L}_{\text{EDSR}}=||\text{EDSR}(N_{\mathcal{T}};\Theta_{2})-\mathbf{B}_{g}||_{2}^{2}.$ (10)
With the training of the SR network completed, we freeze its parameters and apply it to the denoised spike frame $\mathbf{L}_{\text{bsn}}(t)$, yielding the resolution-enhanced spike frame $\mathbf{L}_{\text{bsn}}^{\uparrow}(t)$.
3.3.2 Knowledge Distillation Framework
While the aforementioned processing pipeline achieves the multi-modality fusion of the blurry input and the spike stream, several aspects still need refinement: